A vehicle of mass 550 kg moves at a speed of 16.5 m.s−1 on a circular path of radius 71.0 m. Calculate the magnitude of the centripetal force required to keep the vehicle on this trajectory in N and round to the nearest whole number of Newtons.



1
Expert's answer
2021-04-23T07:28:43-0400

Answer

Centripetal force

"F=\frac{mv^2}{r}"

"F=\frac{550(16.5)^2}{71}=2109.0N"